OpenStudy (anonymous):

integrate 1/sqrt(x^2 +16) dx

OpenStudy (adamaero):

Trig substitution?

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ \sqrt{x^{2} + 16}} dx \]

OpenStudy (anonymous):

yeah, i let x = 4tan solved for dx = sec^2

OpenStudy (anonymous):

im just not getting the right answer.

OpenStudy (adamaero):